OpenStudy (anonymous):

Find dy/dx if y=sin^2 x

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ 1 }{ 2 }-\frac{ 1 }{ 2} \cos2x =\frac{ 1 }{ 2}x+\sin2x\]

OpenStudy (anonymous):

That's because \[\cos2x=1-2\sin ^{2}x \] you get \[\sin^{2}x \] from this relation.

OpenStudy (anonymous):

\[2 \sin(x)\cos(x)\]by the chain rule.